OpenStudy (anonymous):

May I have some help here please, I got a bit confused. There are two triangles one has a perimeter of 14 (sides are y, y, and base x) and the other has a perimeter of 21(sides are 5/4y , 5/4y and base 3x). What are x and y?

OpenStudy (stacey):

Write two perimeter equations and then you will have two equations and two unknowns that you can solve for. The first perimeter equation will be x+2y=14.

OpenStudy (anonymous):

I got that Stacy, thanks for your help! I know the second triangle's perimeter is : 5/4y +5/4y +3x =21 Could you help me some more here?

OpenStudy (anonymous):

2y + x = 14 5/2y + 3x = 21 2y + x = 14 -2(5/2y + 3x = 21) 6(2y + x = 14) -5y - 6x = -42 12y + 6x = 84 -5y - 6x = -42 7y = 42 y = 6 Plug it into an equation. 2(6) + x = 14 12 + x = 14 x = 2 Therefore: x = 2 y = 6 (2, 6)

OpenStudy (anonymous):

@Calcmathlete, thanks a lot for your explanation, but how did you get 5/2, shouldn't it be 5/4?

OpenStudy (anonymous):

I simplified. 5/4 + 5/4 = 10/4 = 5/2

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

ok, I tried that and erased it thinking I was wrong. Thanks man! Yes but not the negative.

OpenStudy (anonymous):

Oh ok.

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

Why do you have -5/2 though?

OpenStudy (anonymous):

Where? It says -2(5/2y + 3x = 21). I'm multiplying the equation by -2.

OpenStudy (anonymous):

oh ok! You are a live saver man! I appreciate you taking the time to explain it to me. :)

OpenStudy (anonymous):

You're welcome and no problem. I'm here to assist. :)